# Solving the PTA Data Analysis Problem with a Global Gibbs Scheme

### v1

@@ I Introduction
After decades of observing millisecond-period pulsars (MSPs), the International Pulsar Timing Array (IPTA) found evidence for a gravitational-wave background (GWB) [5, 6, 7]. The findings leading to the announcement in the summer of 2023 came after years of work by the constituents of the IPTA analyzing different and independent PTA data sets, with the North American Nanohertz Observatory for Gravitational Waves (NANOGrav) reporting the highest statistical significance for a GWB signal across its 67 analyzed MSPs. The findings of European Pulsar Timing Array (EPTA), Indian pulsar timing array (InPTA), and Parkes pulsar timing array (PPTA) uncover compatible and complementary evidence for a GWB, albeit with varying degrees of statistical significance [8].
@@
In particular, a truly model-agnostic (or minimally modeled) per-frequency estimation of the low-frequency correlated and uncorrelated noise of each pulsar is not within the capability of the current GWB detection techniques. Since the signature of the inter-pulsar correlations induced by the GWs from some modified theories of gravity are predicted to vary significantly with GW frequency [17, 18], an agnostic per-frequency Bayesian pipeline could be invaluable to tests of beyond-GR gravity efforts using PTAs. Furthermore, as shown in Kaiser et al. [19], resolving source confusion (i.e., disentangling multiple astrophysical and cosmological sources for a GWB) is an extremely difficult task even when attempted on simulated PTA data sets. Perhaps a model-agnostic per-frequency data analysis technique could provide a better chance of succeeding at such attempts.
Additionally, there is an ongoing need to build a robust pipeline to estimate various global PTA-level quantities (e.g., the GWB correlation signature and its power spectral density) from the results of single-pulsar analyses. This need is based on the fact that many pulsars are affected by chromatic (i.e., radio-frequency dependent) noise whose properties are unique to (the line-of-sight path to) individual pulsars and require bespoke noise modeling. The inclusion of such customized models for every affected pulsar in the global PTA-level analysis is practically impossible at present. Hence, there is a need for techniques capable of achieving this goal.
@@
The paper is structured as follows. In §II, we review the PTA noise modeling structure by defining relevant quantities and introducing the most common noise modeling assumptions. In §III, we perform a detailed derivation and further discuss the new multi-pulsar Gibbs sampling method. In §IV, the implementation challenges and the flow of the Gibbs sampling is presented. In §V, we introduce the basics of the post-processing tools we utilize to process the output of Gibbs sampling. In §VI.3, we explore the capabilities of the presented multi-pulsar Gibbs sampling using realistic and theoretical PTA simulated data sets. In §VII, we present how the result of single-pulsar analysis can be used to approximately characterize an underlying GWB. Finally, in §VIII, we present our concluding remarks.
@@ II.1 Single pulsar
such that the mean is $T\bm{b}$, and the white noise covariance $N$ in each pulsar is composed of squared TOA uncertainties and other calibrating factors (e.g., EFAC, EQUAD, ECORR) that are determined prior to any global PTA analysis. [^1] This procedure of fixing parameters in a global PTA analysis that have previously been searched for at the single-pulsar level can be justified by the fact that these white noise parameters are typically not highly covariant with other processes. The parameters $\bm{b}$ are typically not of primary interest as their covariance across many realizations of the noise is the quantity that can reveal information about the power at low-frequencies in each pulsar as well as in inter-pulsar correlated processes like the GWB. Hence, we introduce $B$ as the covariance of $\bm{b}$, and use Bayes’ Theorem to write
@@
with $\Phi$ as the covariance matrix of dimensions $2n_{f}\times 2n_{f}$ on the Fourier coefficients $\bm{a}$. Notice that the $\bm{\infty}$ simply means that we have placed an unbounded improper prior on the linear timing model offsets; this presents no practical issues since the likelihood is strongly peaked in these parameters, and furthermore we are typically only interested in the inverse covariance matrix, such that $B^{-1}=\mathrm{diag}(0,\Phi^{-1})$. In software implementations, the large numerical value of $10^{40}$ is frequently used instead of $\bm{\infty}$ for practical reasons.
@@ III.1 Prior probability
We note several important points here: $(1)$ this prior applies to the matrix as a whole, rather than to its individual elements independently; $(2)$ the log-uniform prior for single-pulsar analyses in Equation 9 is a special case of this prior when $n_{p}=1$ and $\nu=2$; and $(3)$ even though the above prior guarantees invariance under a scaling transformation, it does not guarantee invariance under a logarithmic transformation of $\Phi_{II}$ parameters as Equation 9 does. Philosophically, we intend for our new prior to behave similarly to a Jeffreys prior, in that it should not prefer one parameterization over another. Additionally, the prior should introduce minimal information to the model relative to the data. To achieve this behavior, careful selection of the parameter $\nu$ is crucial. We discuss the details of construction of the prior in the next section.
@@ IV A multi-pulsar Gibbs scheme
Gibbs sampling [35] is a class of Markov Chain Monte Carlo (MCMC) algorithm in which parameters are updated sequentially through draws from their posteriors conditioned on the fixed values of other parameters. Once each parameter is updated, its value is held fixed in the conditional posterior of the next parameter’s update, and so forth. Despite the fact that random samples are drawn from conditional probabilities of each parameter, over many iterations, the distribution of samples converges to the joint marginal posterior probability of all parameters as guaranteed by the principle of detailed balance. In the field of PTAs, Gibbs sampling has been used previously in van Haasteren and Vallisneri [30], Laal et al. [20] and Wang and Taylor [36].
@@ IV.1 Overview of the algorithm
Here we outline a step-by-step implementation of Gibbs sampling for the multi-pulsar analysis of §II. Note that the treatment of the white noise parameters is identical to that of Laal et al. [20]. No analytic conditional distribution can be found for the white noise parameters, so we must perform a short Metropolis-Hastings MCMC exploration to find a fair draw from the white noise conditional posterior during each loop of the Gibbs sampling. Although we could indeed sample white noise parameters in a global PTA analysis, as previously mentioned, we typically fix these to values determined from single-pulsar analyses. Hence in the following, we assume these are fixed by previous analyses, but the generalization of the Gibbs algorithm to sampling all parameters is trivial.
  • Step 1: Make an initial guesses of $\Phi$, denoted by $\Phi_{0}$. This can be achieved practically by guessing an initial spectrum of PSD values for each pulsar to occupy the diagonal of $\Phi$. For the off-diagonal elements— i.e., $\Phi_{IJ}$ where $I\neq J$— the diagonal elements modulated by relevant values of the Hellings and Downs curve are a decent initial guess.
@@
In practice, since we are sampling large numbers of parameters, it is necessary to impose some constraints on certain values. Constraining $\Phi_{II}$ is necessary to avoid exploring parts of the parameter space that have no support under the prior. While Everson and Morris [39] introduce a method of constraining the diagonals of a Wishart distribution, the method is very computationally challenging and cannot be applied to Inverse-Wishart distributions easily. Instead, we use rejection sampling (i.e., rejecting samples outside a certain fixed boundary) as a way of forcing the $\Phi_{II}$ parameters to be confined within $10^{-11}$ and $10^{-4}$ (units of $\text{seconds}^{2}$) for every frequency bin. With this practical constraint imposed during the Gibbs algorithm above, we now refer to the entire routine as Multi-pulsar Gibbs Sampling (MG) from hereon.
The parameter dimensionality tackled by MG is daunting. A real PTA dataset with $100$ pulsars in which $30$ frequency-bins are used to model low-frequency red noise process has $6,000$ modes in $\bm{b}$, and $166,500$ covariance parameters in $\bm{\Phi}$. Without Gibbs sampling, the inclusion of all these parameters would be an even more daunting exercise, even with gradient-based MCMC methods like No-U-Turn Sampling (NUTS) [40]. Current analyses by PTAs get around this obstacle by assuming a parameterized functional form for the elements of $\Phi$ and the inter-pulsar correlation structure, reducing the size of the parameter space significantly. However, our goal here is to produce posterior samples that are as agnostic as feasible and allow such models to be imposed using MG samples in post-processing. We will discuss this point in §V.
@@ IV.2 Implementation details
Figure 1 showcases a map of the computational tasks at each iteration of Gibbs sampling. The figure is self-explanatory: to move from each node to another, the operation specified by the text surrounding the relevant arrow is performed on the quantity that the arrow originates from which results in the quantity that the arrow points towards. To estimate $\bm{a}$ from $\Phi^{-1}$, first the matrix $\Sigma^{-1}$ is constructed by following Equation 14. Then, the upper-Cholesky-decomposition of $\Sigma^{-1}$, denoted by $U$, is calculated and used within two linear-algebra solvers to eventually draw a random set of $\bm{a}$ from the multivariate Gaussian distribution of Equation 13. Note that $Z$ is a set of standard Gaussian random numbers required to sample from a multivariate normal distribution using the well-known affine transformation technique [^7] The affine transformation states that to sample from $N\left(\bm{\mu},\Sigma\right)$, one can sample from $Z=N\left(0,\mathbb{I}\right)$ and find $\bm{\mu}+\sqrt{\Sigma}Z$, where $\sqrt{\Sigma}$ is a square-root factorization of $\Sigma$.. To continue the flow, the scale-matrix $S$ is constructed for each frequency-bin using Equation 16. For the sole purpose of numerical stabilization of $S$, first, the correlation version of $S$, denoted by $\hat{S}$, is calculated [^8] The correlation version of any covariance matrix is found by element-wise division of the covariance matrix by the matrix constructed from the outer-product of the square-root of its diagonals.. Second, for each pulsar and at each frequency-bin, a random number between $10^{-8}$ and $10^{-5}$ is drawn and added to the diagonals of the correlation matrix resulting in the matrix $\hat{S}^{+}$. The reason for performing such perturbations is entirely for numerical stability. Note that using numbers lower than $10^{-8}$ makes $\hat{S}^{+}$ to have a very high condition-number while using numbers larger than $10^{-5}$ changes the value of the scale-matrix along its diagonals by more than $0.00110^{-5}\times 100=$ 0.001 percent. Furthermore, $\hat{S}^{+}$ is used to reconstruct $S$ which is now more numerically stable since it is cured of its low-rank deficiency. Finally, to complete the iteration and obtain $\Phi$ and $\Phi^{-1}$, a Bartlett decomposition is performed using the standard-Wishart random $n_{p}\times n_{p}$ matrix $A$ and the upper-Cholesky-decomposition, denoted by $P$, of the numerically stable matrix $S$. If the diagonals of the estimated $\Phi$ is not bounded by what the prior demands, the $\Phi$ matrix is discarded and another Bartlett decomposition using a completely different $A$ is performed.
@@ IV.3 Implementation challenges
Figure 1: A schematic representation of the computational tasks carried at each iteration of our multi-pulsar Gibbs sampling. The figure is self-explanatory. The quantity that the arrow is being originated from is used within the operation indicated by the text surrounding the arrow to result in the quantity that arrow is pointing towards. The color and shape coding of the blocks is to help distinguish the calculated quantities from each other. The blue rectangles contain the quantities that must be saved before the next iteration. The green circles contain auxiliary quantities that connect the different pieces together. The red diamonds represent the nodes where the numerical instability of various quantities might prevent the sampling iteration from completing successfully. And finally, the gray ellipses are the nodes where random draws from standard normal and Wishart distributions are made.
@@ IV.4 Improving the convergence rate of Gibbs Sampling
Even though samples from empirical posterior distributions such as from our ENDJ are model dependent and therefore not fully representative of the MG posterior distribution, they serve as effective proposal distributions. By accepting/rejecting with the standard Metropolis-Hastings ratio, we are assured to converge to the true MG posterior. The frequency at which the ENDJ draws are considered is set prior to the the start of the Gibbs sampling. In our implementation, the ENDJ draws are less frequent than the Gibbs sampling draws by at least four orders of magnitude.
@@ V.2 Fitting cross-correlations
where the index $i$ is over the set of $\Phi$ posterior samples drawn using MG. The term being subtracted from $\Phi_{IJ}$ is a power-law model of the GWB’s cross-spectrum, in which $\gamma=13/3$ to represent the expected spectral shape given by a population of supermassive black-hole binaries, and $A$ is the GWB’s characteristic strain amplitude at a frequency of $f_{\mathrm{yr}}=1/\mathrm{year}$. The template correlation pattern $\Gamma_{IJ}^{\text{HD}}$ is the Hellings and Downs function, expected for a statistically isotropic distribution of GWB intensity.
We can also adopt a parameterized model of the correlation signature rather than merely assume that it is the Hellings and Downs function. Moreover, we can do this at any given frequency bin $k$, to characterize the frequency-dependent correlation structure of the GW signal. A simple one-parameter model to fit is the so-called generalized transverse (GT) overlap reduction function [43, 44] given by
@@
such that $\Phi_{k;IJ}=\Gamma_{k;IJ}^{\mathrm{GT}}\rho_{k}$, where $I\neq J$, $x_{IJ}=(1-\cos\xi_{IJ})/2$, and $\xi_{IJ}$ is the angular separation between pulsars $I$ and $J$. The parameters describing the power-spectral-density and shape of the correlations at each frequency are $\rho_{k}$ and $\tau_{k}$, respectively. The GT model includes the Hellings and Downs function as a special case where $\tau=-1$. Figure 2 shows the shape of the GT function for different values of $\tau$ drawn uniformly between $-1.5$ and $1.5$.
@@
Figure 2: The generalized transverse (GT) parametrization of inter-pulsar correlations featured in Equation 28. The dashed lines denote the envelope of correlation curves given by $\tau\in U[-1.5,1.5]$, where solid colored curves correspond to particular random draws of $\tau$. The more positive the value of $\tau$, the closer the resulting pattern is to the top dashed bound. The Hellings and Downs function is a particular case of this family of correlations corresponding to $\tau=-1$.

### v2

@@ I Introduction
After decades of observing millisecond-period pulsars, the International Pulsar Timing Array (IPTA) found evidence for a gravitational-wave background (GWB) [5, 6, 7]. The findings leading to the announcement in the summer of 2023 came after years of work by the constituents of the IPTA analyzing different and independent PTA data sets, with the North American Nanohertz Observatory for Gravitational Waves (NANOGrav) reporting the highest statistical significance for a GWB signal across its 67 analyzed millisecond-period pulsars. The findings of European Pulsar Timing Array (EPTA), Indian pulsar timing array (InPTA), and Parkes pulsar timing array (PPTA) uncover compatible and complementary evidence for a GWB, albeit with varying degrees of statistical significance [8].
@@
In particular, a truly model-agnostic (or minimally modeled) per-frequency estimation of the low-frequency correlated and uncorrelated noise of each pulsar is not within the capability of the current GWB detection techniques. Since the signature of the inter-pulsar correlations induced by the GWs from some modified theories of gravity are predicted to vary significantly with GW frequency [17, 18], an agnostic per-frequency Bayesian pipeline could be invaluable to tests of beyond-GR gravity efforts using PTAs.
Additionally, there is an ongoing need to build a robust pipeline to estimate various global PTA-level quantities (e.g., the GWB correlation signature and its power spectral density) from the results of single-pulsar analyses. This need is based on the fact that many pulsars are affected by chromatic (i.e., radio-frequency dependent) noise whose properties are unique to (the line-of-sight path to) individual pulsars and require bespoke noise modeling. The inclusion of such customized models for every affected pulsar in the global PTA-level analysis is practically impossible at present [19, 20, 21]. Hence, there is a need for techniques capable of achieving this goal.
@@
The paper is structured as follows. In §II, we review the PTA noise modeling structure by defining relevant quantities and introducing the most common noise modeling assumptions. In §III, we perform a detailed derivation and further discuss the new multi-pulsar Gibbs sampling method. In §IV, the implementation challenges and the flow of the Gibbs sampling is presented. In §V, we introduce the basics of the post-processing tools we utilize to process the output of Gibbs sampling. In §VI, we explore the capabilities of the presented multi-pulsar Gibbs sampling using realistic and theoretical PTA simulated data sets. In §VII, we present how the result of single-pulsar analysis can be used to approximately characterize an underlying GWB. Finally, in §VIII, we present our concluding remarks.
@@ II.1 Single pulsar
such that the mean is $T\bm{b}$, and the white noise covariance $N$ in each pulsar is composed of squared TOA uncertainties and other calibrating factors (e.g., EFAC, EQUAD, ECORR defined in Arzoumanian et al. [31]) that are determined prior to any global PTA analysis. [^1] This procedure of fixing parameters in a global PTA analysis that have previously been searched for at the single-pulsar level can be justified by the fact that these white noise parameters are typically not highly covariant with other processes. The parameters $\bm{b}$ are typically not of primary interest as their covariance across many realizations of the noise is the quantity that can reveal information about the power at low-frequencies in each pulsar as well as in inter-pulsar correlated processes like the GWB. Hence, we introduce $B$ as the covariance of $\bm{b}$, and use Bayes’ Theorem to write
@@
with $\Phi$ as the covariance matrix of dimensions $2n_{f}\times 2n_{f}$ on the Fourier coefficients $\bm{a}$. Notice that the $\bm{\infty}$ simply means that we have placed an unbounded improper prior on the linear timing model offsets; this presents no practical issues since the likelihood is strongly peaked in these parameters, and furthermore we are typically only interested in the inverse covariance matrix, such that $B^{-1}=\mathrm{diag}(0,\Phi^{-1})$. This is a common practice in the PTA detection efforts [32, 33, 34]. In software implementations, the large numerical value of $10^{40}$ is frequently used instead of $\bm{\infty}$ for practical reasons.
@@ III.1 Prior probability
We note several important points here: $(1)$ this prior applies to the matrix as a whole, rather than to its individual elements independently; $(2)$ the log-uniform prior for single-pulsar analyses in Equation 9 is a special case of this prior when $n_{p}=1$ and $\nu=2$; and $(3)$ even though the above prior guarantees invariance under a scaling transformation, it does not guarantee invariance under a logarithmic transformation of $\Phi_{II}$ parameters as Equation 9 does. Philosophically, we intend for our new prior to behave similarly to a Jeffreys prior  [^5] Jeffreys prior is a popular choice for a non-informative prior derived form the mathematical structure of the likelihood function., in that it should not prefer one parameterization over another. Additionally, the prior should introduce minimal information to the model relative to the data. To achieve this behavior, careful selection of the parameter $\nu$ is crucial. We discuss the details of construction of the prior in the next section.
@@ IV A multi-pulsar Gibbs scheme
Gibbs sampling [39] is a class of MCMC algorithm in which parameters are updated sequentially through draws from their posteriors conditioned on the fixed values of other parameters. Once each parameter is updated, its value is held fixed in the conditional posterior of the next parameter’s update, and so forth. Despite the fact that random samples are drawn from conditional probabilities of each parameter, over many iterations, the distribution of samples converges to the joint marginal posterior probability of all parameters as guaranteed by the principle of detailed balance. In the field of PTAs, Gibbs sampling has been used previously in van Haasteren and Vallisneri [33], Laal et al. [22] and Wang and Taylor [40].
@@ IV.1 Overview of the algorithm
Here we outline a step-by-step implementation of Gibbs sampling for the multi-pulsar analysis of §II. Note that the treatment of the white noise parameters is identical to that of Laal et al. [22]. No analytic conditional distribution can be found for the white noise parameters, so we must perform a short Metropolis-Hastings MCMC exploration to find a fair draw from the white noise conditional posterior during each loop of the Gibbs sampling. Although we could indeed sample white noise parameters in a global PTA analysis, as previously mentioned, we typically fix these to values determined from single-pulsar analyses. Hence in the following, we assume these are fixed by previous analyses, but the generalization of the Gibbs algorithm to sampling all parameters is, in theory, possible through the methodology implemented and tested previously in Laal et al. [22].
  • Step 1: Make an initial guess of $\Phi$, denoted by $\Phi_{0}$. This can be achieved practically by guessing an initial spectrum of PSD values for each pulsar to occupy the diagonal of $\Phi$. For the off-diagonal elements— i.e., $\Phi_{IJ}$ where $I\neq J$— the diagonal elements modulated by relevant values of the Hellings and Downs [41] (HD) curve are an adequate initial guess.
@@
In practice, since we are sampling large numbers of parameters, it is necessary to impose some constraints on certain values. Constraining $\Phi_{II}$ is necessary to avoid exploring parts of the parameter space that have no support under the prior. While Everson and Morris [44] introduce a method of constraining the diagonals of a Wishart distribution, the method is very computationally challenging and cannot be applied to Inverse-Wishart distributions easily as the proposed method involves a layer of iterative computation in which costly matrix products need to be performed. Instead, we use rejection sampling (i.e., rejecting samples outside a certain fixed boundary) as a way of forcing the $\Phi_{II}$ parameters to be confined within $10^{-11}$ and $10^{-4}$ (units of $\text{seconds}^{2}$) for every frequency bin. With this practical constraint imposed during the Gibbs algorithm above, we now refer to the entire routine as Multi-pulsar Gibbs Sampling (MG) from hereon.
The parameter dimensionality tackled by MG is daunting. A real PTA dataset with $100$ pulsars in which $30$ frequency-bins are used to model low-frequency red noise process has $6,000$ modes in $\bm{b}$, and $166,500$ covariance parameters in $\bm{\Phi}$. Without Gibbs sampling, the inclusion of all these parameters would be an even more daunting exercise, even with gradient-based MCMC methods like No-U-Turn Sampling [45]. Current analyses by PTAs get around this obstacle by assuming a parameterized functional form for the elements of $\Phi$ and the inter-pulsar correlation structure, reducing the size of the parameter space significantly. However, our goal here is to produce posterior samples that are as agnostic as feasible and allow such models to be imposed using MG samples in post-processing. We will discuss this point in §V.
@@ IV.2 Implementation details
Figure 1 showcases a map of the computational tasks at each iteration of Gibbs sampling. To move from each node to another, the operation specified by the text surrounding the relevant arrow is performed on the quantity that the arrow originates from which results in the quantity that the arrow points towards. To estimate $\bm{a}$ from $\Phi^{-1}$, first the matrix $\Sigma^{-1}$ is constructed by following Equation 14. Then, the upper-Cholesky-decomposition of $\Sigma^{-1}$, denoted by $U$, is calculated and used within two linear-algebra solvers to eventually draw a random set of $\bm{a}$ from the multivariate Gaussian distribution of Equation 13. Note that $Z$ is a set of standard Gaussian random numbers required to sample from a multivariate normal distribution using the well-known affine transformation technique [^8] The affine transformation states that to sample from $N\left(\bm{\mu},\Sigma\right)$, one can sample from $Z=N\left(0,\mathbb{I}\right)$ and find $\bm{\mu}+\sqrt{\Sigma}Z$, where $\sqrt{\Sigma}$ is a square-root factorization of $\Sigma$.. To continue the flow, the scale-matrix $S$ is constructed for each frequency-bin using Equation 16. For the sole purpose of numerical stabilization of $S$, first, the correlation version of $S$, denoted by $\hat{S}$, is calculated [^9] The correlation version of any covariance matrix is found by element-wise division of the covariance matrix by the matrix constructed from the outer-product of the square-root of its diagonals.. Second, for each pulsar and at each frequency-bin, a random number between $10^{-8}$ and $10^{-5}$ is drawn and added to the diagonals of the correlation matrix resulting in the matrix $\hat{S}^{+}$. The reason for performing such perturbations is entirely for numerical stability. Note that using numbers lower than $10^{-8}$ makes $\hat{S}^{+}$ to have a very high condition-number while using numbers larger than $10^{-5}$ changes the value of the scale-matrix along its diagonals by more than $0.00110^{-5}\times 100=$ 0.001 percent. Furthermore, $\hat{S}^{+}$ is used to reconstruct $S$ which is now more numerically stable since it is cured of its low-rank deficiency. Finally, to complete the iteration and obtain $\Phi$ and $\Phi^{-1}$, a Bartlett decomposition is performed using the standard-Wishart random $n_{p}\times n_{p}$ matrix $A$ and the upper-Cholesky-decomposition, denoted by $P$, of the numerically stable matrix $S$. If the diagonals of the estimated $\Phi$ is not bounded by what the prior demands, the $\Phi$ matrix is discarded and another Bartlett decomposition using a completely different $A$ is performed.
@@ IV.3 Implementation challenges
Figure 1: A schematic representation of the computational tasks carried at each iteration of our multi-pulsar Gibbs sampling. The quantity that the arrow is being originated from is used within the operation indicated by the text surrounding the arrow to result in the quantity that arrow is pointing towards. The color and shape coding of the blocks is to help distinguish the calculated quantities from each other. The blue rectangles contain the quantities that must be saved before the next iteration. The green circles contain auxiliary quantities that connect the different pieces together. The red diamonds represent the nodes where the numerical instability of various quantities might prevent the sampling iteration from completing successfully. And finally, the gray ellipses are the nodes where random draws from standard normal and Wishart distributions are made.
@@ IV.4 Improving the convergence rate of Gibbs Sampling
Even though samples from empirical posterior distributions such as from our ENDJ are model dependent and therefore not fully representative of the MG posterior distribution, they serve as effective proposal distributions. By accepting/rejecting with the standard Metropolis-Hastings ratio, we are assured to converge to the true MG posterior. The frequency at which the ENDJ draws are considered is set prior to the the start of the Gibbs sampling. In our implementation, the ENDJ draws are less frequent than the Gibbs sampling draws by at least two to four orders of magnitude.
@@ V.2 Fitting cross-correlations
where the index $i$ is over the set of $\Phi$ posterior samples drawn using MG. The term being subtracted from $\Phi_{IJ}$ is a power-law model of the GWB’s cross-spectrum, in which $\gamma=13/3$ to represent the expected spectral shape given by a population of supermassive black-hole binaries, and $A$ is the GWB’s characteristic strain amplitude at a frequency of $f_{\mathrm{yr}}=1/\mathrm{year}$. The template correlation pattern $\Gamma_{IJ}^{\text{HD}}$ is the HD curve, expected for a statistically isotropic distribution of GWB intensity. To distinguish estimates of $A_{\text{GWB}}$ obtained from minimizing Equation 27 with respect to $A$ from those obtained through a Bayesian scheme using Equation 26, we use a “hat” on $A_{\text{GWB}}$ as in $\hat{A}_{\text{GWB}}$.
We can also adopt a parameterized model of the correlation signature rather than merely assume that it is the HD curve. Moreover, we can do this at any given frequency bin $k$, to characterize the frequency-dependent correlation structure of the GW signal. A simple one-parameter model to fit is the so-called generalized transverse (GT) overlap reduction function [48, 49] given by
@@
such that $\Phi_{k;IJ}=\Gamma_{k;IJ}^{\mathrm{GT}}\rho_{k}$, where $I\neq J$, $x_{IJ}=(1-\cos\xi_{IJ})/2$, and $\xi_{IJ}$ is the angular separation between pulsars $I$ and $J$. The parameters describing the power-spectral-density and shape of the correlations at each frequency are $\rho_{k}$ and $\tau_{k}$, respectively. The GT model includes the HD curve as a special case where $\tau=-1$. Figure 2 shows the shape of the GT function for different values of $\tau$ drawn uniformly between $-1.5$ and $1.5$.
@@
Figure 2: The generalized transverse (GT) parametrization of inter-pulsar correlations featured in Equation 28. The dashed lines denote the envelope of correlation curves given by $\tau\in U[-1.5,1.5]$, where solid colored curves correspond to particular random draws of $\tau$. The more positive the value of $\tau$, the closer the resulting pattern is to the top dashed bound. The HD curve is a particular case of this family of correlations corresponding to $\tau=-1$.